1. ∫sin^43xcos^23xdx
2. ∫x^5 sqrt of 1-x^3 dx
1
Expert's answer
2020-09-01T18:20:26-0400

"\u222bsin^4(3x)cos^2(3x)dx=\u222bsin^2(3x)* (sin^2(3x)*cos^2(3x))dx=\u222b(1\/2)*(1-cos(6x))*(1\/4)*sin^2(6x) dx=\u222b(1\/8)*(1-cos(6x))*(1\/2)*(1-cos(12x))dx=\u222b(1\/16)*(1-cos(6x)-cos(12x)+ cos(6x)*cos(12x))dx=\u222b(1\/16)*(1-cos(6x)-cos(12x)+ cos(6x)*cos(12x))dx=\u222b(1\/16)*(1-cos(6x)-cos(12x)+ (1\/2)*cos(6x)+(1\/2)*cos(18x))dx=\u222b(1\/16)*(1-(1\/2)*cos(6x)-cos(12x)+(1\/2)*cos(18x))dx=(1\/16)*(x-(1\/12)*sin(6x)-(1\/12)*sin(12x)+(1\/36)*sin(18x))+C=x\/16-(1\/192)*sin(6x)-(1\/192)*sin(12x)+(1\/576)*sin(18x)+C""\u222bx^5* (1-x^3)^(1\/2) dx=\u222bx^3*x^2*(1-x^3)^(1\/2) dx=\u222b(1-u)*u^(1\/2) d(u\/(-3))=-1\/3\u222bu^(1\/2)-u^(3\/2) du=-1\/3(u^(3\/2)\/(3\/2)-u^(5\/2)\/(5\/2))+C=-(2\/9)*u^(3\/2)+(2\/15)*u^(5\/2)+C=-(2\/9)* (1-x^3)^(3\/2)+(2\/15)*(1-x^3)^(5\/2)+C"